# Huppert’s conjecture for finite simple exceptional groups of Lie type

### v1

@@ 3. Character degrees of simple exceptional groups of Lie type
We are now ready to obtain the results on the character degree sets of finite simple exceptional groups of Lie type. From now on we assume that $q=p^{f}$ is a power of a prime $p,$ where $f\geq 1$ and $H$ is isomorphic to one of the following simple groups:
$\textrm{F}_{4}(q),~{}{}^{2}\textrm{E}_{6}(q),~{}\textrm{E}_{6}(q),~{}\textrm{E}_{7}(q),~{}\textrm{E}_{8}(q),$ (1)
We first begin with the family $\textrm{F}_{4}(q).$ As the case when $H\cong\textrm{F}_{4}(2)$ has been handled in [29], we always assume that $q>3$ when $H\cong\textrm{F}_{4}(q).$
Lemma 3.1.
  1. $(i)$
  2. $(ii)$
  3. $(iii)$
  4. $(iv)$
  5. $(v)$
  6. $(vi)$
    The $p$-part of $x$ is at most $q^{16}$ when $q$ is odd and $\frac{1}{2}q^{13}$ when $q$ is even.
  7. $(vii)$
  8. $(viii)$
  9. $(ix)$
    $H$ has no consecutive degrees.
Proof.
Statements $(i)-(vii)$ can be checked directly given the character degree set of $H$ in [16]. We only need to verify statements $(viii)$ and $(ix).$ For $(viii),$ assume that $x,y\in{\mathrm{cd}}(H)-\{1,q^{24}\}$ with $(x,y)=1.$ By $(i),$ we have that either $(\ell_{1}\ell_{2},x)>1$ or $(\ell_{1}\ell_{2},y)>1.$ Without loss of generality, we assume that $(\ell_{1}\ell_{2},x)>1.$
Assume first that $\ell_{1}\mid x.$ Then $(\ell_{1},y)=1.$ By $(ii),$ either $y=\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{8}$ or $p\mid y.$ If the first case holds, then $\Phi_{3}\Phi_{4}\Phi_{6}\Phi_{8}\mid y$ and so as $(x,y)=1$ we deduce that $(\Phi_{3}\Phi_{4}\Phi_{6}\Phi_{8},x)=1,$ which is impossible by $(iv).$ Thus $p\mid y$ and hence $p\nmid x.$ If $(\ell_{2},y)=1,$ then $y$ is coprime to both $\ell_{1}\ell_{2}$ so $y=\frac{1}{4}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{6}^{2}$ by $(i).$ But then $\Phi_{3}\Phi_{6}\mid y$ and so $(\Phi_{3}\Phi_{6},x)=1.$ Moreover as $p\nmid x,$ we deduce that $(p\Phi_{3}\Phi_{6},x)=1,$ which is impossible by $(iv).$ Thus $\ell_{2}\mid y$ and hence $(\ell_{2},x)=1,$ so $(p\ell_{2},x)=1.$ It follows from $(iii)$ that $x=\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{12}$ or $q$ is odd and $x=\Phi_{3}\Phi_{6}\Phi_{12}.$ If the former case holds then $\Phi_{3}\Phi_{4}\Phi_{6}\mid x$ and so $(\Phi_{3}\Phi_{4}\Phi_{6},y)=1,$ which is impossible by $(iv).$ Therefore $q$ is odd and $x=\Phi_{3}\Phi_{6}\Phi_{12}.$ Using $(iv)$ again, the only possibility for $y$ is $\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}.$ If $p=3,$ then we are done. Now assume that $p\neq 3.$ We will show that $(\Phi_{3}\Phi_{6}\Phi_{12},\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8})>1.$ As $q$ is coprime to $3,$ we deduce that $\Phi_{1}^{4}\Phi_{2}^{4}=(q^{2}-1)^{4}$ is divisible by $3^{4}$ and thus $\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}$ is divisible by $3$ Moreover $\Phi_{3}\Phi_{6}\Phi_{12}=q^{8}+q^{4}+1=(q^{8}-1)+(q^{4}-1)+3$ is also divisible by $3,$ therefore the greatest common divisor of $\Phi_{3}\Phi_{6}\Phi_{12}$ and $\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}$ is divisible by $3$ so it is nontrivial.
Assume that $\ell_{2}\mid x.$ Then $(\ell_{2},y)=1.$ If $(\ell_{1},y)=1,$ then $y=\frac{1}{4}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{6}^{2}$ by $(i).$ Hence $p\Phi_{3}\Phi_{6}\mid y$ so $(p\Phi_{3}\Phi_{6},x)=1,$ which is impossible by $(iv).$ Thus $\ell_{1}\mid y$ and so $(\ell_{1},x)=1.$ By $(ii),$ we obtain that $x=\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{8}$ or $p\mid x.$ If the former case holds, then $\Phi_{3}\Phi_{4}\Phi_{6}\Phi_{8}\mid x$ and so $(\Phi_{3}\Phi_{4}\Phi_{6}\Phi_{8},y)=1,$ which is impossible by $(iv).$ Thus $p\mid x$ and hence $p\nmid y.$ Thus $(p\ell_{2},y)=1.$ It follows from $(iii)$ that $y=\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{12}$ or $q$ is odd and $y=\Phi_{3}\Phi_{6}\Phi_{12}.$ Using the same argument as in the previous case, we obtain the required result.
We next show that $H$ has no consecutive degrees. Assume that $x=y+1.$ By $(vii),$ we can assume that $x>y>1.$ Since $q^{24}\pm 1$ do not divide the order of $H,$ we deduce that $H$ has no such degrees. So, we also assume that $x,y\in{\mathrm{cd}}(H)-\{1,q^{24}\}.$ Moreover $(x,y)=1$ and so by $(viii),$ we must have that $p=3$ and $\{x,y\}=\{\Phi_{3}\Phi_{6}\Phi_{12},\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}\}.$ Observe that $\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}>q^{14}>q^{9}>\Phi_{3}\Phi_{6}\Phi_{12}$ as $q\geq 3,$ therefore $x=\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}~{}\mbox{and}~{}y=\Phi_{3}\Phi_{6}\Phi_{12}.$ But then $y+1<q^{9}+1<q^{14}<x,$ a contradiction. The proof is now complete. ∎
For the remaining lemmas, the proofs are basically identical to that of Lemma 3.1, so we will skip all the detail.
@@ 4. Verifying Step 1: $G^{\prime}=G^{\prime\prime}$
In this section, we will show that $G^{\prime}=G^{\prime\prime}.$ By way of contradiction, suppose that $G^{\prime}\neq G^{\prime\prime}.$ Then there exists a normal subgroup $N\unlhd G$ of $G$ such that $G/N$ is solvable minimal with respect to being nonabelian. By [25, Lemma 2.3], $G/N$ is an $r$-group for some prime $r$ or $G/N$ is a Frobenius group.
@@
Recall that if $\chi\in{\mathrm{Irr}}(G)$, then we say that $\chi$ is isolated in $G$ if $\chi(1)$ is divisible by no proper nontrivial character degree of $G,$ and no proper multiple of $\chi(1)$ is a character degree of $G.$ In this situation, we also say that $\chi(1)$ is an isolated degree of $G.$
@@
Case $1$ $G/N$ is an $r$-group. By [18, Theorem $1.1$], the only nontrivial prime power degree of $H$ is the degree of the Steinberg character of $H,$ which is $|H|_{p}.$ Since $G/N$ is a nonabelian $r$-group, it possesses an irreducible character $\tau$ of degree $r^{b}$ so $\tau(1)=r^{b}=|H|_{p}$ and hence $r=p.$ By Thompson’s Theorem ([13, Corollary 12.2]), $H$ has a nontrivial irreducible character $\psi$ with $p\nmid\psi(1),$ so by [9, Lemma $2$], we have that $\psi_{N}\in{\mathrm{Irr}}(N)$ and hence by Gallagher’s Theorem, $\tau\psi$ is an irreducible character of $G$ of degree $\tau(1)\psi(1)=|H|_{p}\psi(1).$ But this contradicts [25, Lemma 2.4].
@@
Case $2$ $G/N$ is a Frobenius group with kernel $F/N,$ $|F/N|=r^{a},$ $1<m=|G:F|\in{\mathrm{cd}}(G)$ and $r^{a}\equiv 1(\mbox{mod $m$}).$ By [25, Lemma 2.4], we have that ${\mathrm{St}}_{H}(1)=|H|_{p}$ is an isolated degree of $H$ so by [25, Lemma 2.3(b)(2)], either $m=|H|_{p}$ or $r=p.$ We consider each family of simple groups separately.
@@
Subcase $H\cong{\rm F}_{4}(q).$ Recall that $q>2$ in this case. Assume first that $m=|H|_{p}.$ As $r\nmid m,$ we have that $r\neq p.$ By [4, $13.9$], $H$ possesses a unipotent character $\varphi$ labeled by the symbol $\phi_{2,4}^{\prime}$ with degree $\varphi(1)=\frac{1}{2}q\Phi_{4}\Phi_{8}\Phi_{12}.$ As no proper multiple of $m$ is a character of $G$ and $\varphi(1)\nmid m,$ we deduce from [25, Lemma 2.3(b)(1)] that $r^{a}\mid\varphi(1)^{2}.$ As $r\neq p,$ we obtain that $r^{a}\mid\varphi(1)_{p^{\prime}}^{2}=\frac{1}{4}\Phi_{4}^{2}\Phi_{8}^{2}\Phi_{12}^{2}.$ We have $\frac{1}{2}\Phi_{4}\Phi_{8}\Phi_{12}=\frac{1}{2}(q^{4}+1)(q^{6}+1)<q^{10}.$ Hence $r^{a}\leq\varphi(1)_{p^{\prime}}^{2}<q^{20}.$ Since $m\mid r^{a}-1,$ $m=q^{24}\leq r^{a}-1<q^{20}-1<q^{20},$ a contradiction.
@@
Thus $r=p.$ Hence $m\in{\mathrm{cd}}(G)$ is coprime to $p$ and so $m\neq|H|_{p}.$ Let $\chi_{i}\in{\mathrm{Irr}}(H),i=1,2,$ be irreducible characters of $H$ obtained from Lemma 3.1$(v)$. For each $i,$ as no proper multiple of $\chi_{i}(1)$ is a degree of $H,$ [25, Lemma 2.3(b)(2)] implies that $m$ divides $\chi_{i}(1)$ and hence $m$ divide $\gcd(\chi_{1}(1),\chi_{2}(1)),$ which implies that $m$ is coprime to $\ell_{1}\ell_{2}.$ By Lemma 3.1$(i)$, $m=\frac{1}{4}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{6}^{2},$ which is impossible as $m$ is coprime to $p.$
@@
Subcase $H\cong{}^{2}{\rm E}_{6}(q).$ Assume first that $m=|H|_{p}.$ As $r\nmid m,$ $r\neq p.$ By Lemma 3.2$(vii)$, $H$ possesses a unipotent character $\varphi$ labeled by the symbol $\phi_{2,4}^{\prime}$ with degree $\varphi(1)=q\Phi_{8}\Phi_{18}.$ As no proper multiple of $m$ is a character of $G$ and $\varphi(1)\nmid m,$ we deduce from [25, Lemma 2.3(b)(1)] that $r^{a}\mid\varphi(1)^{2}.$ As $r\neq p,$ we obtain that $r^{a}\mid\varphi(1)_{p^{\prime}}^{2}=\Phi_{8}^{2}\Phi_{18}^{2}.$ We have $\Phi_{8}\Phi_{18}=q^{10}-q^{7}+q^{6}+q^{4}-q^{3}+1<q^{15}.$ As $r^{a}\mid\varphi(1)_{p^{\prime}}^{2},$ we deduce that $r^{a}\leq\varphi(1)_{p^{\prime}}^{2}<q^{30}.$ As $m\mid r^{a}-1,$ $m=q^{36}\leq r^{a}-1<q^{30}-1<q^{30},$ which is impossible.
@@
Thus $r=p.$ Then $m\in{\mathrm{cd}}(G)-\{1,|H|_{p}\}$ is coprime to $p.$ By [25, Lemma 2.3(b)(2)] and Lemma 3.2$(v),$ $m$ divides both $\chi_{i}(1),i=1,2.$ As $\chi_{i}(1)$ is coprime to $\ell_{i}$ for $i=1,2,$ we deduce that $m$ is coprime to $\ell_{1}\ell_{2}$ so by Lemma 3.2$(i)$, $m=\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10}$, which is impossible as $p\nmid m.$
The argument for the remaining cases are similar.

### v3

@@ 3. Character degrees of simple exceptional groups of Lie type
We are now ready to list the results on the character degree sets of finite simple exceptional groups of Lie type. From now on we assume that $q=p^{f}$ is a power of a prime $p,$ where $f\geq 1$ and $H$ is isomorphic to one of the following simple groups:
$\textrm{F}_{4}(q),~{}{}^{2}\textrm{E}_{6}(q),~{}\textrm{E}_{6}(q),~{}\textrm{E}_{7}(q),~{}\textrm{E}_{8}(q).$ (1)
Since $H\cong\textrm{F}_{4}(2)$ has been handled in [33], we assume that $q\geq 3$ when $H\cong\textrm{F}_{4}(q).$ As a demonstration, we give a detailed proof for the case $H\cong{}^{2}{\rm E}_{6}(q)$. The remaining cases are similar and we skip their proofs.
@@ 10. $(x)$
Proof.
The lists of character degrees of ${}^{2}{\rm E}_{6}(q)_{sc}$ and ${}^{2}{\rm E}_{6}(q)_{ad}$ are available in [20]. Statements $(i)-(vi)$ can be checked easily. For $(vii)$, see [19]. For $(viii),$ suppose that $x,y\in{\mathrm{cd}}(L)-\{1,q^{36}\}$ and $(x,y)=1.$ Assume first that $(\ell_{1}\ell_{2},x)=1.$ By $(i),$ $x=\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10}$ and so $p\ell_{3}\ell_{4}\mid x.$ Therefore, $(p\ell_{3}\ell_{4},y)=1.$ It follows from $(iv)$ that $q>2$ and $y=\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18}.$ It suffices to show that $p=3.$ Assume that $p\neq 3.$ Then $q^{2}-1$ is divisible by $3$ Now $\Phi_{1}^{4}\Phi_{2}^{4}=(q^{2}-1)^{4}$ is divisible by $3^{4}$ so $x=\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10}$ is divisible by $3$ Also
$\Phi_{3}\Phi_{6}=q^{4}+q^{2}+1=(q^{4}-1)+(q^{2}-1)+3=(q^{2}-1)(q^{2}+2)+3$
is divisible by $3$ and thus $3$ divides both $x$ and $y$ and hence $(x,y)>1,$ a contradiction. Thus $p=3$ as required. Hence we can assume that $(\ell_{1}\ell_{2},x)>1.$ By interchanging the roles of $x$ and $y,$ we can also assume that $(\ell_{1}\ell_{2},y)>1.$ Without loss of generality, we assume that $\ell_{2}\mid x.$ So, $(\ell_{2},y)=1$ hence $\ell_{1}\mid y$ and thus $(\ell_{1},x)=1.$ By $(iii),$ $\ell_{3}\mid x$ so $(\ell_{3},y)=1$ and hence $(\ell_{2}\ell_{3},y)=1.$ By $(ii)$, $p\ell_{1}\ell_{4}\mid y$ and thus $(p\ell_{1}\ell_{4},x)=1,$ which contradicts $(iii)$. This proves claim $(viii)$.
For $(ix),$ without loss of generality, assume that $x=y+1.$ By $(vii),$ we can assume that $x>y>1.$ But $q^{36}\pm 1$ do not divide $|L|,$ so $L$ has no such degree; therefore, we assume that $x,y\in{\mathrm{cd}}(L)-\{1,q^{36}\}.$ Moreover $(x,y)=1$ and so by $(viii),$ $p=3$ and
$\{x,y\}=\{\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10},\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18}\}.$
Since $q\geq 3,$
$\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10}>q^{25}>q^{17}>\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18}$
and thus $x=\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10}~{}\mbox{and}~{}y=\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18}.$ But then
$y+1<q^{17}+1<q^{18}<q^{25}<x$
which is a contradiction as $x=y+1.$
For $(x),$ assume that $13\mid q+$ 1 . Using [20], ${}^{2}{\rm E}_{6}(q)_{sc}$ possesses an irreducible character $\chi$ with degree and multiplicity as in $(x).$ Also $\chi(1)$ divides no other degree of $L.$ Since ${\mathrm{cd}}(H)\subseteq{\mathrm{cd}}(L),$ $\chi(1)$ divides no degree of $H-\{\chi(1)\}.$ Hence it suffices to show that $\chi(1)$ is not a degree of $H.$ Note that $H$ is a normal subgroup of index $3$ in ${}^{2}{\rm E}_{6}(q)_{ad}$ and so, for every irreducible character $\varphi$ of $H,$ either $\varphi(1)$ or $3\varphi(1)$ is a degree of ${}^{2}{\rm E}_{6}(q)_{ad}$ depending on whether $\varphi$ is extendible to ${}^{2}{\rm E}_{6}(q)_{ad}$ or not. Therefore, to show that $\chi(1)$ is not a degree of $H,$ we only need to check that neither $\chi(1)$ nor $3\chi(1)$ is a degree of ${}^{2}{\rm E}_{6}(q)_{ad}.$ Using [20], we check that $\chi(1)$ satisfies these properties, which completes the proof of our first claim. The second claim follows easily using [20]. The proof is now complete. ∎
Lemma 3.2.
  1. $(i)$
  2. $(ii)$
  3. $(iii)$
  4. $(iv)$
  5. $(v)$
  6. $(vi)$
    The $p$-part of $x$ is at most $q^{16}$ when $q$ is odd and at most $\frac{1}{2}q^{13}$ when $q$ is even.
  7. $(vii)$
  8. $(viii)$
  9. $(ix)$
    $H$ has no consecutive degrees.
@@ 4. Verifying Step 1: $G^{\prime}=G^{\prime\prime}$
We verify Step $1$ in this section. Let $\chi$ be an irreducible character of a finite group $G$. Following [29], $\chi$ is isolated in $G$ if $\chi(1)$ is divisible by no proper nontrivial character degree of $G,$ and no proper multiple of $\chi(1)$ is a character degree of $G.$ In this situation, $\chi(1)$ is an isolated degree of $G.$
Proposition 4.1.
Let $G$ be a finite group and let $H$ be one of the finite simple groups in (1). Assume that ${\mathrm{cd}}(G)={\mathrm{cd}}(H)$. Then $G^{\prime}=G^{\prime\prime}.$
Proof.
By way of contradiction, suppose that $G^{\prime}\neq G^{\prime\prime}.$ Then there exists a normal subgroup $N$ of $G$ such that $G/N$ is solvable and minimal with respect to being nonabelian. By [29, Lemma 2.3], $G/N$ is an $r$-group for some prime $r$ or $G/N$ is a Frobenius group.
@@
Case $1$ $G/N$ is an $r$-group. Since $G/N$ is a nonabelian $r$-group, it has an irreducible character $\tau$ of degree $r^{b}$ for some integer $b\geq 1$. By Lemma 2.4, the only nontrivial prime power degree of $H$ is the degree of the Steinberg character of $H,$ which is $|H|_{p}.$ It follows that $\tau(1)=r^{b}=|H|_{p}$ and hence $r=p.$ By Thompson’s Theorem ([16, Corollary 12.2]), $H$ has a nontrivial irreducible character $\psi$ with $p\nmid\psi(1)$. By [12, Lemma $2$], $\psi_{N}\in{\mathrm{Irr}}(N)$ and hence by Gallagher’s Theorem ([16, Corollary 6.17]), $\tau\psi$ is an irreducible character of $G$ of degree $\tau(1)\psi(1)=|H|_{p}\psi(1).$ But this contradicts [29, Lemma 2.4].
@@
Case $2$ $G/N$ is a Frobenius group with kernel $F/N,$ $|F/N|=r^{a},$ $1<m=|G:F|\in{\mathrm{cd}}(G)$ and $r^{a}\equiv 1(\mbox{mod $m$}).$ By [29, Lemma 2.4], ${\mathrm{St}}_{H}(1)=|H|_{p}$ is an isolated degree of $H$ so by [29, Lemma 2.3(b)(2)], either $m=|H|_{p}$ or $r=p.$ We consider each family of simple groups separately.
@@
Subcase $H\cong{\rm F}_{4}(q).$ Recall that $q>2$ in this case. Assume first that $m=|H|_{p}.$ As $r\nmid m,$ $r\neq p.$ By [4, $13.9$], $H$ possesses a unipotent character $\varphi$ labeled by the symbol $\phi_{2,4}^{\prime}$ with degree $\varphi(1)=\frac{1}{2}q\Phi_{4}\Phi_{8}\Phi_{12}.$ As no proper multiple of $m$ is a character of $G$ and $\varphi(1)\nmid m,$ we deduce from [29, Lemma 2.3(b)(1)] that $r^{a}\mid\varphi(1)^{2}.$ As $r\neq p,$ $r^{a}\mid\varphi(1)_{p^{\prime}}^{2}=\frac{1}{4}\Phi_{4}^{2}\Phi_{8}^{2}\Phi_{12}^{2}.$ Now
$\frac{1}{2}\Phi_{4}\Phi_{8}\Phi_{12}=\frac{1}{2}(q^{4}+1)(q^{6}+1)<q^{10}.$
Hence $r^{a}\leq\varphi(1)_{p^{\prime}}^{2}<q^{20}.$ Since $m\mid r^{a}-1,$ $m=q^{24}\leq r^{a}-1<q^{20}-1<q^{20},$ a contradiction.
@@
Thus $r=p.$ Hence $m\in{\mathrm{cd}}(G)$ is coprime to $p$ and so $m\neq|H|_{p}.$ Let $\chi_{i}\in{\mathrm{Irr}}(H),i=1,2,$ be irreducible characters of $H$ obtained from Lemma 3.2$(v)$. For each $i,$ as no proper multiple of $\chi_{i}(1)$ is a degree of $H,$ [29, Lemma 2.3(b)(2)] implies that $m$ divides $\chi_{i}(1)$ and hence $m$ divides $\gcd(\chi_{1}(1),\chi_{2}(1)).$ Thus $m$ is coprime to $\ell_{1}\ell_{2}.$ By Lemma 3.2$(i)$, $m=\frac{1}{4}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{6}^{2},$ which is impossible as $m$ is coprime to $p.$
@@
Subcase $H\cong{}^{2}{\rm E}_{6}(q).$ Assume first that $m=|H|_{p}.$ As $r\nmid m,$ $r\neq p.$ By Lemma 3.1$(vii)$, $H$ possesses a unipotent character $\varphi$ labeled by the symbol $\phi_{2,4}^{\prime}$ with degree $\varphi(1)=q\Phi_{8}\Phi_{18}.$ As no proper multiple of $m$ is a character of $G$ and $\varphi(1)\nmid m,$ by [29, Lemma 2.3(b)(1)], $r^{a}\mid\varphi(1)^{2}.$ As $r\neq p,$ $r^{a}\mid\varphi(1)_{p^{\prime}}^{2}=\Phi_{8}^{2}\Phi_{18}^{2}.$ Now
$\Phi_{8}\Phi_{18}=q^{10}-q^{7}+q^{6}+q^{4}-q^{3}+1<q^{15}.$
As $r^{a}\mid\varphi(1)_{p^{\prime}}^{2},$ $r^{a}\leq\varphi(1)_{p^{\prime}}^{2}<q^{30}.$ Since $m\mid r^{a}-1,$
$m=q^{36}\leq r^{a}-1<q^{30}-1<q^{30},$
which is impossible.
@@
Thus $r=p.$ Then $m\in{\mathrm{cd}}(G)-\{1,|H|_{p}\}$ is coprime to $p.$ By [29, Lemma 2.3(b)(2)] and Lemma 3.1$(v),$ $m$ divides both $\chi_{i}(1),i=1,2.$ As $\chi_{i}(1)$ is coprime to $\ell_{i}$ for $i=1,2,$ $m$ is coprime to $\ell_{1}\ell_{2}$. By Lemma 3.1$(i)$, $m=\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10}$, which is impossible as $p\nmid m.$
The arguments for the remaining cases are similar. ∎